# Elastic scattering at $\sqrt{s}=6$ GeV up to $\sqrt{s}=13$ TeV

### v1

@@ II.3 Non-linear slope
The analysis [73] of the high precision data obtained at $Sp\bar{P}S$ at $\sqrt{s}=541$ GeV in the UA4/2 experiment shows the existence in the slope of the term is proportional $\sqrt{t}$. This term can be related with the nearest $\pi$ meson threshold or, as was shown in [73], such behavior of the differential cross sections can reflect the presence of the contribution of the spin-flip amplitude. As was noted in [79]), the analytic $S-$matrix theory, perturbative quantum chromodynamics and the data require Regge trajectories to be nonlinear complex functions [80, 81]. The Pomeron trajectory has threshold singularities, the lowest one being due to the two-pion exchange required by the $t-$channel unitarity [49]. This threshold singularity appears in different forms in various models (see [79]). In the recent high energy general structure model (HEGS) [39], a small additional term is introduced into the slope which reflects some possible small non-linear properties of the intercept. As a result, the slope is taken in the form
$B(s,t)\ =-\alpha_{1}\ln(\hat{s})t(1-d_{1}t/\ln(\hat{s}e^{d_{2}\alpha_{1}t\ln(\hat{s})}).$ (16)
@@ III Experimental database
In our researches we use widest possible region of experimental data. The energy region begins from $\sqrt{s}=3.5-3.8$ GeV for proton-antiproton scattering. At these energies new data were obtained in the high precision experiment on elastic $p\bar{p}$ scattering at small angles. It has four sets at different energies in the momentum transfer region $|t|=0.000986--0.02$ GeV. The experiment is very important as they obtained the value of $\rho(s,t=0)$ with a remarkably small error and with the size near zero. It is essentially different from the value of $\rho(s,t=0)$ obtained in the framework of the dispersion relation analysis by P. Kroll, which was carried out on the basis of old experimental data with large errors.
Low-energy proton-antiproton data from $\sqrt{s}=11.54$ GeV up to the final ISR energy $\sqrt{s}=62$ GeV are represented in eleven sets. Then we include seven sets of experimental data obtained at the $Sp\bar{p}S$ collider at energies around $\sqrt{s}=540-630$ GeV. The data Tevatron at $\sqrt{s}=1800-1960$ GeV are represented in four sets. The latter data are presented the maximal energy obtained for the proton-antiproton scattering at accelerators. In whole for the $p\bar{p}$ elastic scattering we have 33 sets of the different experiments.
For the proton-proton elastic scattering we take into account the sixty five sets of the different experiments from the low energy $\sqrt{s}=6.1$ GeV up to maximal the LHC energy $\sqrt{s}=13$ TeV. Especially note the high precisions experimental data obtained at small momentum transfer in the FNAL collaborations at $\sqrt{s}=9.8,9.9,10.6,12.3$ GeV and at $\sqrt{s}=19.4,22.2,23.9,27.4$ GeV which started from a very small momentum transfer $t=-0.00049$ GeV2. They is can be compared with the data obtained by the UA4/2 Collaboration at $Sp\bar{p}S$ collider, which is start from $t=-0.000875$ GeV2 and with the data obtained at the LHC by the TOTEM Collaboration at $\sqrt{s}=13$ TeV which reached the $t=-0.00029$ GeV2. Of course, the last case achieved smallest possible angles of scattering. We take into account the experimental data at a high value of momentum transfer up to momentum transfer $-t=10-14$ GeV2. The corresponding experimental data were obtained at energies $\sqrt{s}=19.4,27.4,52.8$ GeV.
@@
We included the data for the spin correlation parameter $A_{N}(s,t)$ of the polarized proton-proton elastic scattering. This set of data includes 235 experimental data for relatively small energies $\sqrt{s}=3.63$ GeV and up to $\sqrt{s}=23.4$ GeV. During our fitting procedure sufficiently good descriptions were obtained.
@@
IV Main amplitudes in the high energy generalized structure (HEGS) model
@@ IV.1 Asymptotic part of the scattering amplitude
The model is based on the idea that at high energies a hadron interaction in the non-perturbative regime is determined by the reggenized-gluon exchange. The cross-even part of this amplitude can have two non-perturbative parts, possible standard pomeron - $(P_{2np})$ and cross-even part of 3-non-perturbative gluons ($P_{3np}$). The interaction of these two objects is proportional to two different form factors of the hadron. This is the main assumption of the model. The second important assumption is that we choose the slope of the second term four times smaller than the slope of the first term, by analogy with the two pomeron cuts. Both terms have the same intercept.
@@ IV.2 Form factors
However, it is not obvious that strong interactions have to be proportional to the electromagnetic properties of hadrons. Taking into account this fact in one of the famous models, Bourrely-Soffer-Wu used some modification of the form factor with free parameters which was obtained from the description of the differential cross section of the hadron scattering. For example, it takes the form
@@
However, this allows some freedom in the $t$-dependence of the scattering amplitude. It is necessary to take into account the parton distribution (PDFs) in the hadrons. But PDFs depend on the Bjorken variable $x$. In the 80s some relations between PDfs and hadron form factor there were proposed.
Note that the function like GPDs(x,t, $\xi=0$) was used already in the old ”Valon” model proposed by Sanevich and Valin in 1986 [23]. In the model, the hadron elastic form factor was obtained by the integration function $L(x)G(x,t)$ where $L(x)$ corresponds to the parton function and $G(x)$ corresponds to the additional function which depends from momentum transfer and $x$. In modern language $L(x)G(x,t)$ exactly corresponds to the GPDs. The scattering amplitude will be
@@
. This form factor can be obtained taking into account PDF of particle interactions which is multiplied by some function depending on momentum transfer $t$ and Boirken variable $x$.
Many different forms of the $t$-dependence of GPDs were proposed. In the quark di-quark model the form of GPDs consists of three parts - PDFs, function distribution and Regge-like function. In other works (see e.g. [59]), the description of the $t$-dependence of GPDs was developed in a more complicated picture using the polynomial forms with respect to $x$.
Commonly, the form of $GPDs(x,\xi,t)$ is determined through the exclusive deep inelastic processes of type $\gamma^{*}p\rightarrow Vp,$ where $V$ stands for a photon or a vector meson. However, such processes have a narrow region of momentum transfer and in most models the $t$-dependence of GPDs is taken in factorization form with the Gaussian form of the $t$-dependence. Really, this form of $GPDs(x,\xi,t)$ can not be used to build space structure of the hadrons, as for that one needs to integrate over $t$ in a maximally wide region.
The conjunction between the momentum transfer and the impact parameter allows one to obtain a space parton distribution that has a probability conditions [14]. The connections between the deep-inelastic scattering, from which we can obtain the $x$-dependence of parton distributions, and the elastic electron-nucleon scattering, where the form factors of the nucleons are obtained, can be derived by using the sum rules [19, 20, 21, 22]. The form factors, which are obtained in different reactions, can be calculated as the Mellin moments of GPDs. Using the electromagnetic (calculated as the zero Mellin moment of GPDs) and gravitomagnetic form factors (calculated as the first moment of GPDs) in the hadron scattering amplitude, one can obtain a quantitative description of hadron elastic scattering in a wide region of energy and transfer momenta.
Using the electromagnetic (calculated as the zero Mellin moment of GPDs) and gravitomagnetic form factors (calculated as the first moment of GPDs) in the hadron scattering amplitude one can obtain a quantitative description of hadron elastic scattering in a wide region of energy and transfer momenta.
@@
where $e_{u}=2/3$ and $e_{d}=-1/3$ are the relevant quark electric charges. As a result, the $t$-dependence of the $GPDs(x,\xi=0,t)$ can be determined from the analysis of the nucleon form factors for which experimental data exist in a wide region of momentum transfer. It is a unique situation as it unites elastic and inelastic processes.
In the limit $t\rightarrow 0$, the functions $H^{q}(x,t)$ reduce to usual quark densities in the proton:
@@
normalized to the number of $u$ and $d$ valence quarks in the proton. The energy-momentum tensor $T_{\mu\nu}$ [24, 20, 21, 13] contains three gravitation form factors (GFF) $A^{Q,G}(t)$, $B^{Q,G}(t)$, and $C^{Q.G}(t)$. We will scrutinize the first one which corresponds to the matter distribution in a nucleon. This form factor contains quark and gluon contributions
@@
To obtain the true form of the proton and neutron form factors, it is important to have the true form of the momentum transfer dependence of GPDs. Let us choose the $t$-dependence of GPDs in a simple form ${\cal{H}}^{q}(x,t)\ =q(x)\ exp[a_{+}\ f(x)\ t],$ with $f(x)=(1-x)^{2}/x^{\beta}$ [99]. The isotopic invariance can be used to relate the proton and neutron GPDs; hence we have the same parameters for the proton and neutron GPDs.
@@
The hadron form factors will be obtained by integration over $x$ in the whole range of $x$ $(0-1)$. Hence the obtained form factors will be dependent on the $x$-dependence of the forms of PDF at the ends of the integration region. The Collaborations determined the PDF sets from the inelastic processes only in some region of $x$, which is only approximated to $x=0$ and $x=1$. Some PDFs have the polynomial form of $x$ with different power. Some other have the exponential dependence of $x$. As a result, the behavior of PDFs, when $x\rightarrow 0$ or $x\rightarrow 1$, can impact the form of the calculated form factors.
@@
$\tilde{G}_{d}(t)=1/(1+q/a_{1}+q^{2}/a_{2}^{2}+q^{3}/a_{3}^{3})^{2}$
which is slightly different from the standard dipole form on two additional terms with small sizes of coefficients. The matter form factor
[…]
@@
Now let us calculate the moments of the GPDs with inverse power of $x$. It gives us the Compton form factors. The results of our calculations of the Compton form factors coincide well with the existing experimental data. $R_{V}(t)$ and $R_{T}(t)$ have a similar momentum transfer dependence but differ essentially in size. On the contrary, the axial form factor $R_{A}$ has an essentially different $t$ dependence.
A good description of the variable form factors and elastic scattering of hadrons gives a large support of our determination of the momentum transfer dependence of GPDs. Based on this determination of GPDs, one can calculate the gravitomagnetic radius of the nucleon using the integral representation of the form factor and make the numerical differentiation over $t$ as $t\rightarrow 0$. This method allows us to obtain a concrete form of the form factor by fitting the result of the integration of the GPDs over $x$. As a result, the gravitomagnetic radius is determined as
@@
One can obtain the gravitational form factors of quarks which are related to the second moments of GPDs. For $\xi=0$, one has
@@
The parameters of the phenomenological form of GPDs can be obtained from the analysis of the experimental data for the proton and neutron electromagnetic form factors simultaneously. Our determination of the momentum transfer dependence of GPDs of hadrons allows us to obtain good quantitative descriptions of different form factors, including the Compton, electromagnetic, transition and gravitomagnetic form factor simultaneously.
V Model approximation
@@
$A_{N}\frac{s^{2}}{4\pi}\frac{d\sigma}{dt}=$ (30)
$-[Im(\Phi_{1}(s,t)+\Phi_{2}(s,t)+\Phi_{3}(s,t)-\Phi_{4})(s,t)\Phi^{*}_{5}(s,t)]$
The HEGS model [38, 39] takes into account all five spiral electromagnetic amplitudes. The electromagnetic amplitude can be calculated in the framework of QED. In the high energy approximation, it can be obtained for the spin-non-flip amplitudes:
$\Phi^{em}_{1}(t)=\alpha f_{1}^{2}(t)\frac{s-2m^{2}}{t};\Phi^{em}_{3}(t)=\Phi^{em}_{1};$ (31)
and for the spin-flip amplitudes, with the electromagnetic and hadronic interactions included, every amplitude $\Phi_{i}(s,t)$ can be described as
@@
where $\varphi(s,t)=\varphi_{C}(t)-\varphi_{Ch}(s,t)$, and $\varphi_{C}(t)$ will be calculated in the second Born approximation in order to allow the evaluation of the Coulomb-hadron interference term $\varphi_{Ch}(s,t)$. The quantity $\varphi(s,t)$ has been calculated at large momentum transfer including the region of the diffraction minimum [77, 78].

### v3

@@ II.3 Non-linear slope
The analysis [29] of the high precision data obtained at $SP\bar{P}S$ at $\sqrt{s}=541$ GeV in the UA4/2 experiment shows the existence in the slope of the term proportional to $\sqrt{|t|}$. This term can be related with the nearest $\pi$ meson threshold or, as was shown in [29], such behavior of the differential cross sections can reflect the presence of the contribution of the spin-flip amplitude. As was noted in [56]), the analytic $S-$matrix theory, perturbative quantum chromodynamics and the data require Regge trajectories to be nonlinear complex functions [57, 58]. The Pomeron trajectory has threshold singularities, the lowest one being due to the two-pion exchange required by the $t-$channel unitarity [51]. This threshold singularity appears in different forms in various models (see [56]). In the recent high energy general structure model (HEGS) [34], a small additional term is introduced into the slope which reflects some possible small nonlinear properties of the intercept. As a result, the slope is taken in the form
$B(s,t)\ =-\alpha^{\prime}\ln(\hat{s})[1-d_{1}t/\ln(\hat{s})e^{d_{2}\alpha_{1}t\ln(\hat{s})}].$ (15)
@@ III Experimental database
In our research, we use the widest region of experimental data. The energy region begins from $\sqrt{s}=3.5-3.8$ GeV for proton-antiproton scattering. At these energies new data were obtained in the high precision experiment on elastic $p\bar{p}$ scattering at small angles. It has four sets at different energies in the momentum transfer region $|t|=[0.000986-0.02]$ GeV [59]. This experiment is very important as they obtained the value of $\rho(s,t=0)$ with a remarkably small error and with the size near zero. It is essentially different from the value of $\rho(s,t=0)$ obtained in the framework of the dispersion relation analysis by P. Kroll, which was carried out on the basis of old experimental data with large errors.
Low-energy proton-antiproton data from $\sqrt{s}=11.54$ GeV up to the final ISR energy $\sqrt{s}=62$ GeV are represented in eleven sets [61]. Then we include seven sets of experimental data obtained at the $SP\bar{P}S$ collider at energies around $\sqrt{s}=540-630$ GeV. The Tevatron data at $\sqrt{s}=1800-1960$ GeV are represented in four sets. The latter data present the maximal energy obtained for the proton-antiproton scattering at accelerators. On the whole, for the $p\bar{p}$ elastic scattering we have 33 sets of the different experiments.
For the proton-proton elastic scattering we take into account sixty five sets of different experiments from the low energy $\sqrt{s}=6.1$ GeV up to the maximal LHC energy $\sqrt{s}=13$ TeV [62, 63]. We especially note the high precision experimental data obtained at a small momentum transfer by the FNAL collaborations at $\sqrt{s}=9.8,9.9,10.6,12.3$ GeV and at $\sqrt{s}=19.4,22.2,23.9,27.4$ GeV which start from a very small momentum transfer $t=-0.00049$ GeV2. They can be compared with the data obtained by the UA4/2 Collaboration at the $SP\bar{P}S$ collider, which start from $t=-0.000875$ GeV2 and with the data obtained at the LHC by the TOTEM Collaboration at $\sqrt{s}=13$ TeV which reached the $t=-0.00029$ GeV2. Of course, the last case achieved the smallest possible angles of scattering. We take into account the experimental data at a high value of momentum transfer up to momentum transfer $-t=10-14$ GeV2. The corresponding experimental data were obtained at energies $\sqrt{s}=19.4,27.4,52.8$ GeV.
@@
We included in our simultaneous analysis the data for the spin correlation parameter $A_{N}(s,t)$ of the polarized proton-proton elastic scattering. This set of data includes 235 experimental data for relatively small energies $\sqrt{s}=3.63$ GeV and up to $\sqrt{s}=23.4$ GeV. During our fitting procedure sufficiently good descriptions of the experimental data were obtained.
@@
IV Main amplitudes of the high energy generalized structure (HEGS) model
@@ IV.1 Asymptotic part of the scattering amplitude
The model is based on the idea that at high energies a hadron interaction in the non-perturbative regime is determined by the non-perturbative Regge gluons exchange. The cross-even part of this amplitude can have two non-perturbative parts, possible standard non-perturbative QCD pomeron - $(P_{2np})$ and cross-even part of 3-non-perturbative gluons ($P_{3np}$). The interaction of these two objects is proportional to two different form factors of the hadron and both form factors are discussed in the next subsection. This is the main assumption of the model. Both terms have the same intercept. This corresponds to the maximal Odderon, which was introduced by L. Lukazsuk, B. Nicolescu [64].
Note, the Odderon that arises naturally in perturbative QCD, for example [65], with the intercept $\alpha_{Odd}(0)=1$. They show that shadowing corrections decreased Odderon contributions. However, the experimental data show the Odderon contribution at ISR energy $\sqrt{s}=52.8$ GeV and at Tevatron energy [66] and probably at LHC energy $\sqrt{s}=13$ TeV. For example, in [67] it was noted ”On the other hand, it is possible to introduce the Odderon phenomenologically as an object which does not violate first principles and the axiomatic theorems. In fact it was stated in [28] that the new TOTEM result is a definitive confirmation of the experimental discovery of the Odderon in its maximal form”.
In our fitting procedure we checked up such a possibility and made the fit with two intercepts. As a result we obtained the Odderon intercept $\alpha_{0-Odd}-1=0.1101\pm 0.0004$ (with fixed Pomeron intercept $\alpha_{0-Pom}-1=0.11$). and that $\chi^{2}$ did not practically changes. Hence, this confirms our assumption about the equality of both intercepts.
The second important assumption is that we choose the slope of the second term four times smaller than the slope of the first term, by analogy with the two pomeron cuts. An estimation of the diffraction slope of the generalized BFKL pomeron was made in [68] in terms of the correlation radius for the perturbative gluons. They show that indeed the pomeron trajectory has the finite slope $\alpha^{\prime}_{IP}\sim R^{2}_{c}$ , where $R_{c}$ is the correlation radius for the perturbative gluons. They note that in the gBFKL dynamics a dimensionful $\alpha^{\prime}_{IP}$ is a noneperturbative quantity related to the nonperturbative infrared parameter of the model - the gluon propagation radius $Rc$. For the preferred $Rc=0.27$ fm, quite a small $\alpha^{\prime}_{IP}=0.072$ GeV?2 is found. In our work we used the phenomenological properties of the Pomeron (Odderon) taking into account that the Odderon only feels the centre of the proton and not the pion cloud. Therefore, it is reasonable to assume that the Odderon slope, $B_{Odd}$, is lower than that for the even-signature (Pomeron) amplitude [67]. This takes into account the data of the ISR experiments which show that the slope of the differential cross sections after the second bump is approximately four times smaller than the slope at small $t$ as well as the conclusion of P. Landshoff that the Odderon contributions are essentially important at large momentum transfer. Our fitting procedure confirms also such assumption when we take the second slope as a free parameter.
@@ IV.2 Form factors
Figure 4: The Born amplitudes of nucleon-nucleon elastic scattering: (top) with two non-perturbative gluons - $P_{2np}$ and (bottom) three non-perturbative gluons - $P_{3np}$.
@@
However, it is not obvious that strong interactions have to be proportional to the electromagnetic properties of hadrons. Taking into account this fact in one of the famous models, Bourrely-Soffer-Wu [69] used some modification of the form factor with free parameters which was obtained from the description of the differential cross section of hadron scattering. For example, it takes the form
@@
However, this allows some freedom in the $t$-dependence of the scattering amplitude. It is necessary to take into account the parton distribution (PDFs) in the hadrons. However, PDFs depend on the Bjorken variable $x$. In the 80s some relations between PDFs and hadron form factor were proposed.
Note that the function like GPDs(x,t, $\xi=0$) was used already in the old ”Valon” model proposed by Sanevich and Valin in 1986 [70]. In the model, the hadron elastic form factor was obtained by the integration function $L(x)G(x,t)$ where $L(x)$ corresponds to the parton function and $G(x,t)$ corresponds to the additional function which depends on momentum transfer and $x$. In modern language, $L(x)G(x,t)$ exactly corresponds to the GPDs. The scattering amplitude is
@@
. This form factor can be obtained taking into account PDF of interacting particles which is multiplied by some function depending on momentum transfer $t$ and Boirken variable $x$.
Many different forms of the $t$-dependence of GPDs were proposed [71]. In the quark di-quark model [72] the form of GPDs consists of three parts - PDFs, function distribution and Regge-like function. In other works (see e.g. [73]), the description of the $t$-dependence of GPDs was developed in a more complicated picture using the polynomial forms with respect to $x$.
Commonly, the form of $GPDs(x,\xi,t)$ is determined through the exclusive deep inelastic processes of type $\gamma^{*}p\rightarrow Vp,$ where $V$ stands for a photon or a vector meson. However, such processes have a narrow region of momentum transfer and in most models the $t$-dependence of GPDs is taken in factorization form with the Gaussian form of the $t$-dependence. Really, this form of $GPDs(x,\xi,t)$ can not be used to build a space structure of hadrons, as for that one needs to integrate over $t$ in a maximally wide region.
The conjunction between the momentum transfer and the impact parameter allows one to obtain a space parton distribution that has probability conditions [11]. The connections between the deep-inelastic scattering, from which we can obtain the $x$-dependence of parton distributions, and the elastic electron-nucleon scattering, where the form factors of the nucleons are obtained, can be derived by using the sum rules [17, 18, 19, 20]. The form factors, which are obtained in different reactions, can be calculated as the Mellin moments of GPDs. Using the electromagnetic (calculated as the zero Mellin moment of GPDs) and gravitomagnetic form factors (calculated as the first moment of GPDs) in the hadron scattering amplitude, one can obtain a quantitative description of hadron elastic scattering in a wide region of energy and transfer momenta.
@@
where $e_{u}=2/3$ and $e_{d}=-1/3$ are the relevant quark electric charges. As a result, the $t$-dependence of the $GPDs(x,\xi=0,t)$ can be determined from the analysis of the nucleon form factors for which experimental data exist in a wide region of momentum transfer. This is a unique situation as it unifies elastic and inelastic processes.
In the limit $t\rightarrow 0$, the functions $H^{q}(x,t)$ are reduced to usual quark densities in the proton:
@@
normalized to the number of $u$ and $d$ valence quarks in the proton. The energy-momentum tensor $T_{\mu\nu}$ [74, 18, 19, 21] contains three gravitation form factors (GFF) $A^{Q,G}(t)$, $B^{Q,G}(t)$, and $C^{Q.G}(t)$. We will scrutinize the first one which corresponds to the matter distribution in a nucleon. This form factor contains the quark and gluon contributions
@@
To obtain the true form of the proton and neutron form factors, it is important to have the true form of the momentum transfer dependence of GPDs. Let us choose the $t$-dependence of GPDs in a simple form ${\cal{H}}^{q}(x,t)\ =q(x)\ exp[a_{+}\ f(x)\ t],$ with $f(x)=(1-x)^{2}/x^{\beta}$ [75]. The isotopic invariance can be used to relate the proton and neutron GPDs; hence, we have the same parameters for the proton and neutron GPDs.
@@
The hadron form factors will be obtained by integration over $x$ in the whole range of $x$ $(0-1)$. Hence, the obtained form factors depend on the $x$-dependence of the forms of PDF at the ends of the integration region. The Collaborations determined the PDF sets from the inelastic processes only in some region of $x$, which is only approximated to $x=0$ and $x=1$. Some PDFs have the polynomial form of $x$ with different power. Others have the exponential dependence of $x$. As a result, the behavior of PDFs, when $x\rightarrow 0$ or $x\rightarrow 1$, can impact the form of the calculated form factors.
@@
$\tilde{G}_{d}(t)=1/(1+a_{1}q+q^{2}/a_{2}^{2}+a_{3}^{3}q^{3})^{2}$
which is slightly different from the standard dipole form by two additional terms with small sizes of coefficients ($a_{1}=0.06$, GeV-1, $a^{2}_{2}=0.78$ GeV2, and $a^{3}_{3}=0.08$ GeV-3 ). The matter form factor
$A(t)=$ $\int^{1}_{0}x\ dx[q_{u}(x)e^{2\alpha_{H}f(x)_{u}/t}$
$+q_{d}(x)e^{2\alpha_{H}f_{d}(x)/t}]$
@@
Now let us calculate the moments of the GPDs with inverse power of $x$. This gives us the Compton form factors. The results of our calculations of the Compton form factors coincide well with the existing experimental data. $R_{V}(t)$ and $R_{T}(t)$ have a similar momentum transfer dependence but differ essentially in size. On the contrary, the axial form factor $R_{A}$ has an essentially different $t$ dependence.
A good description of the variable form factors and elastic scattering of hadrons gives large support to our determination of the momentum transfer dependence of GPDs. Based on this determination of GPDs, one can calculate the gravitomagnetic radius of the nucleon using the integral representation of the form factor and make the numerical differentiation over $t$ as $t\rightarrow 0$. This method allows us to obtain a concrete form of the form factor by fitting the result of the integration of the GPDs over $x$. As a result, the gravitomagnetic radius is determined as
@@
One can obtain gravitational form factors of quarks, which are related to the second moments of GPDs. For $\xi=0$, one has
@@
The parameters of the phenomenological form of GPDs can be obtained from the analysis of the experimental data for the proton and neutron electromagnetic form factors simultaneously. Our determination of the momentum transfer dependence of GPDs of hadrons allows us to obtain good quantitative descriptions of different form factors, including the Compton, electromagnetic, transition, and gravitomagnetic form factor simultaneously.
V Model approximation and the structure of the elastic scattering amplitude
@@
$A_{N}\frac{s^{2}}{4\pi}\frac{d\sigma}{dt}=$
$-[Im(\Phi_{1}(s,t)+\Phi_{2}(s,t)+\Phi_{3}(s,t)-\Phi_{4})(s,t)\Phi^{*}_{5}(s,t)].$ (29)
The HEGS model [33, 34] takes into account all five spiral electromagnetic amplitudes. The electromagnetic amplitude can be calculated in the framework of QED. For the spin-flip amplitudes, with the electromagnetic and hadronic interactions included, every amplitude $\Phi_{i}(s,t)$ can be described as
@@
where $\varphi(s,t)=\varphi_{C}(t)-\varphi_{Ch}(s,t)$, and $\varphi_{C}(t)$ are calculated in the second Born approximation in order to allow the evaluation of the Coulomb-hadron interference term $\varphi_{Ch}(s,t)$. The quantity $\varphi(s,t)$ has been calculated at a large momentum transfer including the region of the diffraction minimum [78, 79].
In the standard perturbative picture the spin-flip amplitudes die with growing of energy. However, there are different non-perturbative approaches which show non-dying spin-flip amplitudes, for example, [22, 23, 24, 25] One of these is by E. Kuraev’s [104, 105], which we used for our non-dying part of the spin-flip amplitude. We used also the second part of the spin-flip amplitude which has a Regge form (eq.50) and has the energy dependence of order $1/s$.